OpenStudy (anonymous):

Is the given system of linear equation consistent (_,_), inconsistent or dependent 6x +2y =-4 and 6x +3y =-3

OpenStudy (anonymous):

Any math tutors available

OpenStudy (anonymous):

A system is said to be consistent if it has a unique solution or an infinite number of solutions. A system is inconsistent if it has no solutions.

OpenStudy (anonymous):

The system is consistent and has the solution: x = -1 y = 1

OpenStudy (anonymous):

two linear equations in two variables a1x + b1y + c1 = 0 and a2x + b2y + c2 = 0 a1 b1 will be consistent ( i.e have a unique solution) if ----- ╪ ---- a2 b2 a1 b1 c1 will be inconsistent (i.e. have no solution) if ---- = ----- ╪ ----- a2 b2 c2 a1 b1 c1 will be dependent (i.e. have infinite solutions) if ----- = ----- = ----- a2 b2 c2

OpenStudy (anonymous):

6x +2y =-4 and 6x +3y =-3 In case of yr equations a1 6 ---- = ----- = 1 a2 6 b1 2 ---- = ----- b2 3 c1 4 ---- = ----- c2 3

OpenStudy (anonymous):

since a1 b1 ----- ╪ ---- a2 b2 so it is a consistent pair of equations

OpenStudy (anonymous):

Note that even dependent system is consistent....

OpenStudy (anonymous):

Let me say that it is much less messy if you approach systems of linear equations using linear algebra. You can show that a system of linear equations is equivalent to Ax = b where A is a matrix consisting of the coefficients of the system, x is the unknown and b is the vector in the target space. The problem boils down to the question, is there a vector x such that the image of x under the transformation A is b. If a vector or a set of vectors exist the system is said to be consistent. Row reduction, which is kernel preserving (solution preserving) can be used to find the solution set.

OpenStudy (anonymous):

@Alchemista I beg to differ........☺ if u are moderately good in maths, u can easily calculate the ratios and decide........

OpenStudy (anonymous):

Gaussian elimination is efficient and obviously for larger systems you will be using that approach anyways.

OpenStudy (anonymous):

yes for LARGER systems...........☺